write an equation for the line described: a line whose slope =3 and passes through the point (-2,7)
Answer by AnlytcPhil(1116)

write an equation for the line described: a line whose 
slope =3 and passes through the point (-2,7)


The point-slope formula

y - y1 = m[x - x1]

gives the equation of the line with slope = m which 
passes through the point (x1,y1)

So plug in 3 for m, -2 for x1, 
and 7 for y1:

y - 7 = 3[x - (-2)]

y - 7 = 3[x + 2]

y - 7 = 3x + 6

    y = 3x + 13

Edwin